OpenStudy (anonymous):

For the past 5 months Kaia's gift shop had these profits or losses: –$120, $85, –$60, –$35, and $40. What was her average profit or loss for these five months? A. Her average loss was $68 per month. B. Her average loss was $18 per month. C. Her average profit was $18 per month. D. Her average profit was $68 per month

OpenStudy (anonymous):

cm on gotta leave here in 10 minitues

OpenStudy (anonymous):

will give medal who answers frist

OpenStudy (anonymous_user):

B.

OpenStudy (jack1):

its just "mean" average so add them all together then divide by the total number of inputs

OpenStudy (jack1):

so sum of all = -$90 and total number of inputs is 5 so answer = 90/5 = @anonymous_user 's answer

OpenStudy (anonymous):

A

OpenStudy (anonymous):

who answerd frist

OpenStudy (anonymous_user):

I did

OpenStudy (anonymous):

well you were wreong it was a sorry

OpenStudy (jack1):

@tntrunner500 totally not A buddy, unless you put the wrong profit / loss info in the original question (–$120) + ($85) + (–$60) + (–$35) + ($40) = $125 - $215 = -$90 overall 5 months so divide 90 by 5 = -$18 per month ... so ur answer should be B, and @anonymous_user was right... or the info u wrote in the q was wrong